OpenStudy (anonymous):

Solve the problem. Find all values of k so that the given points are \sqrt {29} units apart. (-5, 5), (k, 0) a. 3, 7 b. 7 c. -3, -7 d. -7

OpenStudy (anonymous):

\[\sqrt{(k+5)^2+5^2}=\sqrt{29}\] is a start

OpenStudy (anonymous):

in other words \[(k+5)^2+25=29\] solve for k

OpenStudy (anonymous):

\[(k+5)^2=4\] \[k+5=2\] or \[k+5=-2\] \[k=-3\] or \[k=-7\]

OpenStudy (anonymous):

Thankyou very much